Solve the equation:
x^-2 - 15x^-1 + 44 = 0
: Solve the equation:
x^-2 - 15x^-1 + 44 = 0

Answer by Alan3354(1444)

x^-2 - 15x^-1 + 44 = 0
Multiply by x^2
1 - 15x + 44x^2 = 0
Solved by pluggable solver: SOLVE quadratic equation (work shown, graph etc)
Quadratic equation ax^2+bx+c=0 (in our case 44x^2+-15x+1 = 0) has the following solutons:

x[12] = (b+-sqrt( b^2-4ac ))/2\a

For these solutions to exist, the discriminant b^2-4ac should not be a negative number.

First, we need to compute the discriminant b^2-4ac: b^2-4ac=(-15)^2-4*44*1=49.

Discriminant d=49 is greater than zero. That means that there are two solutions:  x[12] = (--15+-sqrt( 49 ))/2\a.

x[1] = (-(-15)+sqrt( 49 ))/2\44 = 0.25
x[2] = (-(-15)-sqrt( 49 ))/2\44 = 0.0909090909090909

Quadratic expression 44x^2+-15x+1 can be factored:
44x^2+-15x+1 = (x-0.25)*(x-0.0909090909090909)
Again, the answer is: 0.25, 0.0909090909090909. Here's your graph:
graph( 500, 500, -10, 10, -20, 20, 44*x^2+-15*x+1 )


x = 0.25, x=1/11